UT 20161123 LUCI1 AO Nighttime Commissioning/FLAO-tests

Logs from JPower

Observer LBTO: JPower, JHill
Observer AO: DMiller
Telescope Operator: DHuerta
Mountain support: JMorris

Plan for the night

- AO/LUCI com when conditions suitable

- LN 1⁄2 and pointing model data otherwise

Summary

LN started the night. Seeing was 0.95\x94 on the Dimm so the call was made to switch
to LUCI comm at 19:00. AO/LUCI Comm took over the rest of the night. Seeing ranged from
0.37-0.95\x94 on the DIMM throughout the night, hanging around 0.6\x94 for the majority of it. Targets
Observed include Q104 (Ks only), TTau (BrGam and J), Trapezium (BrGam only, some 400
mode some 300 mode bin 1), and RMon (Complete). Tests with 300 and 400 mode were made
in morning twilight on an AO ref star.

Issues

Note that the Retro reflector brackets were left on. IDL issue early on required a work
around to get NCPA to work. DOug will be contacting Alfio to see if something better can be
worked out. Struggled to close the loop with Bin 1 400 modes but other binning/modes were ok
otherwise. Noticed highly negative countdowns with LUCI. Restarting GEIRS did not resolve.
Once script has a guide star that fell outside the patrol field after the initial offset causing preset
to fail. Had to find new GS.

Details

All times are UT
Files are luci1.20161023.NNNN.fits

01:15 Seeing ~0.95\x94 and holding. Doug Miller has been contacted and is coming in. LN wrapping up. Light on SX side.

01:58 LN handed over for FLAO testing. Authorized LUCI LUCI.

02:18 Sent Preset to Q 0104 Bin4 66 Modes, Ks, #0001-0060 Guider FWHM 0.65-0.88\x94
Readme mentions Ks is sufficient, so moving on to next target which is compatible with FLAO testing

03:56 Preset sent to TTau

04:00 TTau Bin1 400modes BrGamma 30061-0062, No NCPA for
luci1.20161123.0061-62. Critical IDL error preventing NCPA from running. Doug fixed it!

04:20 NCPA back up and running, TTau Bin1 400modes, Restarted back at first position,
#0063-0064 DIMM 0.5\x94, Guider 0.62\x94. Second offset failed at 21:24.

04:28 Resent preset TTau Came back in Bin2, good experiment for whether we should apply NCPA for Bin2 or not.

04:34 Resumed #0065-0066, Then failed offset again.

04:37 Resent preset, Preset failed shortly after sending: Failed Interrupt Status Call.

04:41 Resent preset to TTau, #0067
Science Bin2 with NCPA #0068-0069 FAILED Preset again

04:51 Resent preset to TTau. Preset failed. Also seeing highly neg exposure times (neg 40 seconds).
Restarting GEIRS while Doug troubleshoots

04:59 Geirs restarted. Resent preset. Preset failed again. AO requests not even showing up
in arbitrator. Restarted AO arbitrator.

05:05 Resent preset TTAU BrGamma. Bin2 without NCPA #0070

05:09 Science TTAU BrGamma Bin2 153 modes without NCPA #0071-0094
Seeing Guider 0.37\x94- DIMM 0.65\x94. Successfully offsetting without NCPA. Seeing negative
times up to -48sec.

06:00 Sent preset to Standard HD27560 BrGamma. Forced to Bin2 153 modes to match Science.
DIMM 0.6-0.7\x94 Guider 0.4-0.5\x94. #0095-0107

06:15 Sent preset to Standard HD27560 J. Forced to Bin2 153 modes to match Science.
DIMM 0.6-0.7\x94 Guider 0.4-0.5\x94. #0108-0114

06:26 Sent Preset to TTau J #115

06:31 Science TTAU J Bin2 153 modes without NCPA #0116-0127 Seeing Guider 0.39\x94- DIMM 0.65-79\x94.

06:55 Preset to Trapezium sent and failed. Guide star selected outside patrol field. Editing script. Pointing check.

07:17 Resent preset, New Guide star successful! Blind Blind #0128.

07:23 Trapezium Bin 1 400 modes #0129 Loop opened

07:26 Offset Failed.

07:29 Resending Trapezium preset. #0130

07:35 Science Bin 1 400 modes #0131 Open Loop. Error on Offset on next move.

07:40 Resending Trapezium preset. Blind Blind #0132 Telescope
communication hiccup. Reloading and resending 0:41 Blind Blind #0133

07:46 Trapezium Science Bin 1 400 modes #0134 Open Loop again, Doug doing
some proactive troubleshooting before next offset to see if we can continue.

07:55 Resuming to next offset Bin 1 400 modes Loop closed with NCPA off. Left Blind Filter in
by accident, image #0135-0136 blind.

08:00 Restarted with NCPA back on to see if that was issue with closing loop. Trapezium Bin 1
400 modes #0137-0140 Seeing DIMM 0.58\x94 GUIDER 0.43\x94
luci1.20161123.0141 Loop Opened, offset failed.

08:16 Turned off NCPA. Bin 1 400 modes Starting at point where loop opened in script.
#0142-0146, DIMM 0.56\x94 GUIDER 0.44\x94. Extreme negative countdowns!! -40 seconds.

08:24 Turning back on NCPA Bin 1 400 modes. Tried to resume script after pause but script
disappeared. Reloaded from random point within just to test NCPA on.
#0147-0149 DIMM 0.56\x94 GUIDER 0.44\x94. Open loop #0150,
optical gain problem

08:42 Resending preset, starting Trapezium script from beginning. Troubles getting NCPA back up.

08:56 Resent Trapezium Preset.

09:02 Trapezium Science Bin 1 400 modes NCPA on #0153 Open loop, fail

09:06 Resending preset, starting Trapezium script from beginning. Bin 1 400 modes NCPA OFF
this time #0154

09:12 Trapezium Science Bin 1 400 modes NCPA OFF #0155 Open loop, fail

09:15 Resending preset, starting Trapezium script from beginning. Bin 1 400 modes NCPA ON
this time #0156

09:20 Trapezium Science Bin 1 400 modes NCPA ON #0157 Open loop, fail.

09:28 Tried to resend after recovery and got another error: \x93Illegal Command for state
AOarbitrator. Ready.\x94 More recovery.

09:32 Reset Script: the script starts without hitting go?? Preset didn\x92t send properly.

09:35 Reloaded script and resent preset starting Trapezium script from beginning. Bin 1 400
modes NCPA ON this time #0161-0164, DIMM 0.7\x94 GUIDER 0.51\x94
#0165 Loop Opened,failed. (Negative countdown -57 seconds.) Script failed.

09:51 Resending preset, starting Trapezium script from beginning. Bin 1 300 modes NCPA OFF
this time #0166

09:56 Trapezium Science Bin 1 300 modes NCPA OFF #0167-0186 Completed Br Gam Only.

10:37 Preset to RMon. Bin 3

10:43 Science RMon Bin 3 Seeing varied, DIMM 0.37-1.07\x94 GUIDER 0.43-85\x94 0187-0270

13:02 Sending preset to test AO star, HeI +J crossed Bin 1 300modes

#0271 (0,0)

#0272 (0,5)

#0273 (5,0)

13:14 Sending preset to same test AO star, Bin 1 400modes, failure Needed to do a pointing check,
Failure Failure again, Doug resent preset. Bin 1 400modes

#0274 (0,0)

#0275 (5,0)

#0276 (0,5)

13:25 Performing offsets to close loop (20,0), (0, 20), (-20,0), (0, 40)

13:45 Closing up

13:50 Taking internal flats with N30 for data taken tonight.

N30 Ks Halo2 5 x (3 x 12) #0281-0285 OFF, 0286-0290 ON

N30 z Halo1 5 x (2 x 2.52) #0293-0297 OFF #0298-0302 ON

N30 J Halo2 5 x (4 x 13) #0307-0311 OFF #0312-0016 ON

N30 H Halo2 5 x (3 x 8.5) 0319-0323 OFF #0324-0328 ON

N30 BrGam Halo1 5 x (3 x 5) #0331-0335 OFF #0336-0341 ON

-- %USERSIG{JochenHeidt - 2016-11-23}%
Topic revision: r1 - 23 Nov 2016, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback